Given the equation N2(g)+3H2(g)--->2NH3(g) if ∆H=-92.2kJ and ∆S=-0.1987kJ/K, what is ∆G for reaction at 475°C and at 5°C? Provide the answer in kJ. In each case, is the reaction spontaneous?

1
Expert's answer
2022-04-22T12:10:03-0400

∆G = ∆H - T∆S

1 case: T = 4750C = 748K; ∆G = ∆H - T∆S = -92.2 – 748 * (-0.1987) = 56.43 (kJ) > 0 nonspontaneous.

2 case: T = 50C = 278K; ∆G = ∆H - T∆S = -92.2 – 278 * (-0.1987) = -36.96 (kJ) < 0 spontaneous.